Please explain in-depth how to calculate this and read the tables please!Brief Exercise G-9 Click here to view the factor table 1. Click here to view the factor table 2. Click here to view the factoTable 1 Future Value of 1 (n) Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% 1.00000 1.04000 1.08160 1.12486 1.16986 1.21665 1.265Table 2 Future Value of an Annuity of 1 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% Payments 1.00000 2.04000 3.12160 4.24646 5.41632 6.Table 3 Present Value of 1 5% 6% 7% 8% 9% 10% 11% 12% 15% Periods 96154 95238 92456 90703 .88900 .86384 .85480 .82270 .82193+90 Table 4 Present Value of an Annuity of 1 (n) Payments 4% 5% 96154 95238 1.88609 1.85941 2.77509 2.72325 3.62990 3.54595 4
Answer #1

1) Present Value = amount to be received *( PV at r for n Periods)

Amount =34,100

r=4%

n=9

PV=Present Value at 4% for 9 Periods =0.70259

Answer: Preasent Value = 34,100*0.70259 = $23,958.32(Answer)

2) Present Value = amount to be received *PVA at r for n Periods

Amount =34,100

r=4%

n=6

PV= Present Value at 4% for 9 Periods = 5.24214

Present Value =34,100*5.24214 = $178,756.97 (Answer)